During the Cold War the US and the former Soviet U
Answer #1

Given that :

Diameter of the sphere pellet, D = 4 x 10-3 m

radius of the sphere pellet, r = 2 x 10-3 m

packing fraction of sphere, PFrac = 0.8

Volume of sphere pellet is given by, VP = (4/3) \pir3

VP = (4/3) (3.14) (2 x 10-3 m)3

VP = 3.35 x 10-8 m3

How many liters of shot does Ben need to carry out his demonstration?

Volume of metal is given by,   Vm = VP / (0.8)

Vm = (3.35 x 10-8 m3) / (0.8)

Vm = 4.18 x 10-8 m3

Convert m3 into L :

V = 4.18 x 10-5 Liter